What is the [OH-] of a solution that has a pH of 12.5?


1
Expert's answer
2021-03-29T06:08:21-0400

pOH + pH =14

pOH + 12.5 = 14

pOH = 14 - 12.5

pOH = 1.5

[OH-] = 10-pOH

[OH-] = 10-1.5

[OH-] = 0.0316 M